OpenStudy (anonymous):

1. (4/5)(-3/8) = -12/40, -3/10, or -32/15 2. -2/5(-6/7) = 12/35, 6/7, or -12/35

OpenStudy (anonymous):

\[ \frac{a}{b}\cdot\frac{c}{d}=\frac{a\cdot c}{b \cdot d}\]

OpenStudy (anonymous):

So for \[1) \ \frac{4}{5}\cdot \frac{-3}{8} = \frac{4 \cdot (-3)}{5 \cdot 8}\]